# 1]Meta Superintelligence Labs, Menlo Park, CA 94025, USA

### v1

@@ Abstract
We propose Generalized Primal Averaging (GPA), an extension of Nesterov’s method in its primal averaging formulation that addresses key limitations of recent averaging-based optimizers such as single-worker DiLoCo and Schedule-Free (SF) in the non-distributed setting. These two recent algorithmic approaches improve the performance of base optimizers, such as AdamW, through different iterate averaging strategies. Schedule-Free explicitly maintains a uniform average of past weights, while single-worker DiLoCo performs implicit averaging by periodically aggregating trajectories, called pseudo-gradients, to update the model parameters. However, single-worker DiLoCo’s periodic averaging introduces a two-loop structure, increasing its memory requirements and number of hyperparameters. GPA overcomes these limitations by decoupling the interpolation constant in the primal averaging formulation of Nesterov. This decoupling enables GPA to smoothly average iterates at every step, generalizing and improving upon single-worker DiLoCo. Empirically, GPA consistently outperforms single-worker DiLoCo while removing the two-loop structure, simplifying hyperparameter tuning, and reducing its memory overhead to a single additional buffer. On the Llama-160M model, GPA provides a 24.22% speedup in terms of steps to reach the baseline (AdamW’s) validation loss. Likewise, GPA achieves speedups of 12% and 27% on small and large batch setups, respectively, to attain AdamW’s validation accuracy on the ImageNet ViT workload. Furthermore, we prove that for any base optimizer with regret bounded by $\mathcal{O}(\sqrt{T})$, where $T$ is the number of iterations, GPA can match or exceed the convergence guarantee of the original optimizer, depending on the choice of interpolation constants.
@@ 1 Introduction
As large language models (LLMs) demonstrate increasingly remarkable capabilities at scale (achiam2023gpt; grattafiori2024llama3herdmodels; liu2024deepseek), the pre-training phase has become one of the most expensive stages in the language model training pipeline, often costing hundreds of millions of dollars per run. This significant investment has driven the development of training algorithms and optimizers that enhance the efficiency, scalability, and robustness of language model pre-training. One significant area of research is the design of training algorithms for scalable distributed learning. In this area, the DiLoCo algorithm has emerged as the leading practical approach (diloco; liu2024asynchronous; douillard2025streaming; charles2025communication).
Despite its practical success, the underlying reasons for DiLoCo’s effectiveness remains poorly understood. Importantly, DiLoCo is not limited to distributed training: single-worker DiLoCo outperforms AdamW even in the non-distributed setting. kallusky2025snoo suggest that this is due to its novel combination of the Nesterov optimizer with the Lookahead method (lookahead), also known as Step-$K$ Nesterov. The method first accumulates multiple updates from a base optimizer on an inner set of weights, forming what is called a pseudo-gradient. It then applies Nesterov momentum to these pseudo-gradients to update an outer set of weights, and finally resets the inner weights to match the new values of the outer weights. On a 160 million parameter Llama model, single-worker DiLoCo achieves speedups of up to 20.06% in terms of steps to reach the final validation loss by AdamW; see Figure 1(b).
(a) Both GPA and single-worker DiLoCo, when using AdamW as their base optimizer, outperform the AdamW baseline for training a 160M parameter Llama model. Notably, increasing the number of inner steps (up to 16) improves the performance of single-worker DiLoCo, while GPA updates the parameters at every step. However, GPA uses a heuristic to choose its interpolation constants to match the number of inner steps for single-worker DiLoCo.
(b) We evaluate the speedup achieved by single-worker DiLoCo and GPA in reducing the number of steps required to reach the final validation loss attained by AdamW, across different effective numbers of inner steps. GPA attains the highest speedup of 24.22% when the effective inner steps is equal to 16. Single-worker DiLoCo only outperforms AdamW for 8 and 16 inner steps.
Figure 1: Comparison of validation loss and speedup for AdamW, single-worker DiLoCo, and GPA. Although setting the inner steps = 32 yields a lower final validation loss (see Figure 1(a)), setting the inner steps = 16 is faster in terms of number of steps to attain the target validation loss (see Figure 1(b)).
@@
  • Our experiments demonstrate that GPA consistently outperforms single-worker DiLoCo and AdamW on dense 160 million and 1 billion parameter language models. This is further validated on the ImageNet ViT workloads on both small and large batch settings. In particular, on the Llama-160M model, we find that GPA provides speedups of 24.22% in terms of steps to reach the baseline validation loss. Likewise, GPA obtains speedups of 12% and 27% on small and large batch setups on the ImageNet ViT workload, respectively.
  • We provide a theoretical justification for GPA through convergence guarantees that demonstrate improved convergence over the base optimizer under some circumstances in the stochastic convex and non-smooth setting.
@@ 2.1 Different Formulations of Nesterov Momentum
where $\mu>0$ is the momentum hyperparameter and $b^{(t)}\in\mathbb{R}^{n}$ is the momentum buffer initialized at $b^{(0)}=0$. An alternative formulation, which we call the modern formulation, is used by software libraries such as PyTorch [^2] https://docs.pytorch.org/docs/2.8/generated/torch.optim.SGD.html and JAX [^3] https://optax.readthedocs.io/en/latest/api/optimizers.html#optax.sgd due to its ease of use:
@@
In both formulations, we maintain a momentum buffer that averages the gradients seen throughout the training process. However, unlike Sutskever’s formulation (equation 2), the modern formulation (equation 3) uses the iterate $x^{(t)}$ directly for the gradient computation, rather than the ancillary point $x^{(t)}+\mu b^{(t-1)}$, simplifying its practical implementation. If both formulations are run side-by-side with the same seed, they will evaluate gradients at exactly the same points, but their validation losses at iterates $x^{(t)}$ for each method will differ.
@@
Unlike the Sutskever and modern formulations framed in equations 2 and 3, the primal averaging formulation in equation 4 explicitly names two iterate sequences: a sequence where the gradients (or, more generally, the search directions) are computed at, i.e., the gradient computation sequence $\{y^{(t)}\}_{t=1}^{T}$, as well as another sequence used for model evaluation that accumulates a running average of updated iterates $\{z^{(t)}\}_{t=1}^{T}$, i.e., the model evaluation sequence $\{x^{(t)}\}_{t=1}^{T}$. Since $y^{(t)}$ interpolates the smoothed sequence $x^{(t)}$ and unsmoothed sequence $z^{(t)}$, it increases the contribution of the gradient update to $y^{(t)}$ compared to $x^{(t)}$. This explicit formulation is convenient for implementation and theoretical analysis, and naturally leads to a view of acceleration as built upon iterate averaging, rather than from the physics-inspired intuition of gradient averaging behind momentum that is more commonly introduced.
@@
Given fixed learning rates $\gamma_{\operatorname{primal}},\gamma_{\operatorname{modern}}>0$, the primal averaging formulation of Nesterov’s method (equation 4) is equivalent to its modern formulation (equation 3) in the sense that
$y_{\operatorname{primal}}^{(t)}=x_{\operatorname{modern}}^{(t)}\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ b_{\operatorname{modern}}^{(t)}=\frac{1}{\left(1-\mu\right)\gamma_{\operatorname{primal}}}\left(x_{\operatorname{primal}}^{(t)}-x_{\operatorname{primal}}^{(t+1)}\right),$ (5)
@@
The proof of this simple statement is rather technical, so we defer it to Appendix 10. Similar formulations and equivalences can be derived for Polyak momentum (polyak1964some; defazio2020mom; ziyin2020laprop); see Appendix 8.
@@ 2.2 Single-Worker DiLoCo and its Weaknesses
DiLoCo was originally introduced as a distributed algorithm for cross-datacenter training (diloco). The method computes multiple inner steps of the base optimizer on the inner weights, then applies Nesterov (equation 3) on the average pseudo-gradient, the difference between the previous and updated inner model weights, to update the outer weights. The inner weights are then reset to the outer weights.
DiLoCo requires storing two additional optimizer states of the same shape as the model parameters: the momentum buffer $b^{(t)}$ and the current model parameters $x^{(t)}$ (also known as the outer weights). DiLoCo’s handling of fast inner weights and slow outer weights can be interpreted as a modified Lookahead method that applies Nesterov momentum to the outer weight updates (lookahead). The method was recently analyzed in khaled2025understanding, and demonstrated significant compute factor gains in the non-distributed setting in kallusky2025snoo.
A simplified version of non-distributed or single-worker DiLoCo with $H$ inner steps of the base optimizer can be described as:
$\begin{split}p^{(t)}&=x^{(t)}-\operatorname{\texttt{BaseOptIteration}}(x^{(t)};\{\gamma^{(j)}\}_{j=1}^{H},H)\\ b^{(t)}&=\mu b^{(t-1)}+p^{(t)}\\ x^{(t+1)}&=x^{(t)}-\tilde{\gamma}[\mu b^{(t)}+p^{(t)}],\end{split}$ (6)
@@
Weaknesses in DiLoCo’s hierarchical framework. However, this two-level structure is undesirable. From an algorithmic perspective, one would prefer to average iterates on-the-fly, as opposed to averaging trajectories that implicitly contain multiple iterations of the base optimizer. From the users’ perspective, the two-level structure introduces an additional copy of the model weights required to compute the pseudo-gradient, and introduces additional hyperparameters to tune, e.g., the inner and outer learning rates, momentum, and number of inner steps. Lastly, from the distributed training perspective, DiLoCo couples the number of inner steps as a hyperparameter for both local SGD as well as for the modified Nesterov algorithm, causing the algorithm’s performance to counterintuitively improve as the number of inner steps increases. One would instead expect that communicating more often should always be beneficial. These challenges motivate the development of a new algorithm that removes the two-level structure while offering a separate hyperparameter that can smoothly average the observed iterates at every iteration.
@@ 2.3 Schedule-Free Learning
In parallel, Schedule-Free learning (SF) (schedule-free) was recently proposed as a wrapper to any base optimizer using a variant of the primal averaging formulation of Nesterov’s method (equation 4) for hyperparameter-free learning:
@@
Originally designed to eliminate the need for manually specified learning rate schedules, Schedule-Free has demonstrated the surprising ability to not only match, but even surpass the practical performance of the original base optimizer. This is done by decoupling the momentum hyperparameter used in the $x^{(t)}$ and $y^{(t)}$ sequences, unlike the standard primal averaging formulation of Nesterov (equation 4). Through the choice of $\mu$, the method interpolates between uniform Polyak-Ruppert averaging and stochastic primal averaging (ruppert; polyak; tao2018).
@@ 3 Generalized Primal Averaging (GPA)
By decoupling the constants for the model evaluation and gradient computation sequences in the primal averaging formulation of Nesterov’s method (equation 4) and leveraging the observation of using exponential moving averaging in lieu of uniform averaging in Schedule-Free (equation 7), we introduce the Generalized Primal Averaging (GPA) framework:
@@
Unlike the modern formulation of Nesterov momentum (equation 3) or DiLoCo (equation 6) built on (pseudo-) gradient averaging, GPA is defined based on the primal or iterate averaging framework. We argue that this provides a more meaningful characterization of the method. For example, the primal averaging interpretation naturally extends to other search directions by replacing $g(y^{(t)};\xi^{(t)})$ with the search direction $d^{(t)}$ evaluated at $y^{(t)}$. This extension is not intuitive from the gradient averaging perspective, as it would translate to averaging search directions (with potentially different, evolving preconditioners) in the momentum buffer.
Learning rate schedules. By replacing Polyak-Ruppert averaging with exponential moving averaging, GPA is not inherently schedule-free and requires the use of a learning rate schedule. To see why, observe that Polyak averaging places increasingly less weight $1/(t+1)$ on the most recent iterate $z^{(t+1)}$, which plays a similar role to learning rate scheduling (sandler2023trainingtrajectories; schedule-free). GPA instead places a constant weight $\mu_{x}$ on the most recent iterate $z^{(t+1)}$ by leveraging an exponential moving average, thereby requiring a learning rate schedule compared to Schedule-Free. This is reflected theoretically in their last-iterate convergence properties.
@@
Other properties. GPA also retains several desirable properties of the base optimizer for deep learning. Because $\mu_{x},\mu_{y}\in[0,1]$, GPA preserves modular norm bounds of the model parameters. Additionally, when $\mu_{y}>0$, GPA can be implemented with only one extra copy of the model weights – specifically, by storing $y^{(t)}$ and reconstructing $x^{(t)}$ from $y^{(t)}$ and $z^{(t)}$ during evaluation – unlike DiLoCo, which demands more memory overhead. More details on these properties are provided in Appendix 9.

### v3

# arXiv:2512.17131v3[cs.LG] 27 Feb 2026

@@ Abstract
We propose Generalized Primal Averaging (GPA), an extension of Nesterov’s method that unifies and generalizes recent averaging-based optimizers like single-worker DiLoCo and Schedule-Free, within a non-distributed setting. While DiLoCo relies on a memory-intensive two-loop structure to periodically aggregate pseudo-gradients using Nesterov momentum, GPA eliminates this complexity by decoupling Nesterov’s interpolation constants to enable smooth iterate averaging at every step. Structurally, GPA resembles Schedule-Free but replaces uniform averaging with exponential moving averaging. Empirically, GPA consistently outperforms single-worker DiLoCo and AdamW with reduced memory overhead. GPA achieves speedups of 8.71%, 10.13%, and 9.58% over the AdamW baseline in terms of steps to reach target validation loss for Llama-160M, 1B, and 8B models, respectively. Similarly, on the ImageNet ViT workload, GPA achieves speedups of 7% and 25.5% in the small and large batch settings respectively. Furthermore, we prove that for any base optimizer with $\mathcal{O}(\sqrt{T})$ regret, where $T$ is the number of iterations, GPA matches or exceeds the original convergence guarantees depending on the interpolation constants.
@@ 1 Introduction
As large language models (LLMs) demonstrate increasingly remarkable capabilities at scale (achiam2023gpt; grattafiori2024llama3herdmodels; liu2024deepseek), the pre-training phase has become one of the most resource intensive stages in the language model training pipeline. This has encouraged the development of training algorithms and optimizers that enhance the efficiency, scalability, and robustness of language model pre-training. One significant area of research is the design of training algorithms for scalable distributed learning. In this area, the DiLoCo algorithm has emerged as the leading practical approach (diloco; liu2024asynchronous; douillard2025streaming; charles2025communication).
Despite its practical success, the underlying reasons for DiLoCo’s effectiveness remain poorly understood. Importantly, DiLoCo is not limited to distributed training: single-worker DiLoCo outperforms AdamW even in the non-distributed setting. kallusky2025snoo suggest that this is due to its novel combination of the Nesterov optimizer with the Lookahead method (lookahead), called Step-$K$ Nesterov. The method accumulates multiple updates from a base optimizer on an inner set of weights, forming what is called a pseudo-gradient. It then applies Nesterov momentum to the pseudo-gradient to update an outer set of weights, and subsequently resets the inner weights to match the new values of the outer weights. On a 160 million parameter Llama model, single-worker DiLoCo achieves speedups of up to 6.32% in terms of steps to reach the final validation loss by AdamW; see Figure 1(b).
(a) Both GPA and single-worker DiLoCo, when using AdamW as their base optimizer, outperform the tuned AdamW baseline for training a 160M parameter Llama model. Notably, increasing the number of inner steps (up to 64) improves the performance of single-worker DiLoCo. GPA instead updates the parameters at every step using a heuristic to choose interpolation constants that approximately match the number of inner steps for single-worker DiLoCo.
(b) Speedup achieved by single-worker DiLoCo and GPA measured in terms of reduction in number of steps required to attain the final validation loss obtained by AdamW, across different effective numbers of inner steps. GPA attains the highest speedup of 8.71% when the effective inner steps is equal to 32. Single-worker DiLoCo does not exhibit any speedup for 128 inner steps.
Figure 1: Comparison of validation loss and speedup for AdamW, single-worker DiLoCo, and GPA on Llama-160M model. See Appendix 11.2 for similar results on Llama-1B.
@@
  • Our experiments demonstrate that GPA consistently outperforms single-worker DiLoCo and AdamW on dense 160 million and 1 billion parameter language models. We validate our results on different modalities through an 8 billion parameter Llama code generation model and through a vision workload using ImageNet ViT in both small- and large-batch settings. In particular, on the Llama-160M, 1B, and 8B models, we find that GPA provides speedups of 8.71%, 10.13%, and 9.58%, respectively, in terms of steps to reach the baseline validation loss. Likewise, GPA obtains speedups of 7% and 25.5% (small and large batch) on the ImageNet ViT workloads.
  • We provide theoretical justification for GPA by proving convergence guarantees that show improved convergence over the base optimizer on stochastic convex and non-smooth functions under some settings.
@@ 2.1 Different Formulations of Nesterov Momentum
where $\mu>0$ is the momentum hyperparameter and $b^{(t)}\in\mathbb{R}^{n}$ is the momentum buffer initialized at $b^{(0)}=0$. An alternative formulation, which we call the modern formulation, is used by software libraries such as PyTorch [^2] PyTorch SGD documentation and JAX [^3] Optax (JAX) SGD documentation due to its ease of use:
@@
In both formulations, we maintain a momentum buffer that averages the gradients seen throughout the training process. However, unlike Sutskever’s formulation (Equation˜2), the modern formulation (Equation˜3) uses the iterate $x^{(t)}$ directly for the gradient computation, rather than the ancillary point $x^{(t)}+\mu b^{(t-1)}$, simplifying its practical implementation. If both formulations are run side-by-side with the same seed, they will evaluate gradients at exactly the same points, but their validation losses at iterates $x^{(t)}$ for each method will differ.
@@
Unlike the Sutskever and modern formulations framed in equations 2 and 3, the primal averaging formulation in Equation˜4 explicitly names two iterate sequences: a sequence where the gradients (or, more generally, the search directions) are computed at, i.e., the gradient computation sequence $\{y^{(t)}\}_{t=1}^{T}$, as well as another sequence used for model evaluation that accumulates a running average of updated iterates $\{z^{(t)}\}_{t=1}^{T}$, i.e., the model evaluation sequence $\{x^{(t)}\}_{t=1}^{T}$. Since $y^{(t)}$ interpolates the smoothed sequence $x^{(t)}$ and unsmoothed sequence $z^{(t)}$, it increases the contribution of the gradient update to $y^{(t)}$ compared to $x^{(t)}$. This explicit formulation is convenient for implementation and theoretical analysis, and naturally leads to a view of acceleration as built upon iterate averaging, rather than from the physics-inspired intuition of gradient averaging behind momentum that is more commonly introduced.
@@
Given fixed learning rates $\gamma_{\operatorname{primal}},\gamma_{\operatorname{modern}}>0$, the primal averaging formulation of Nesterov’s method (Equation˜4) is equivalent to its modern formulation (Equation˜3) in the sense that
$y_{\operatorname{primal}}^{(t)}$ $=x_{\operatorname{modern}}^{(t)}\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}$ (5)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ b_{\operatorname{modern}}^{(t)}$ $=\frac{1}{\left(1-\mu\right)\gamma_{\operatorname{primal}}}\left(x_{\operatorname{primal}}^{(t)}-x_{\operatorname{primal}}^{(t+1)}\right),$
@@
The proof of this simple statement is rather technical, so we defer it to Appendix 10. Similar formulations and equivalences can be derived for Polyak momentum (polyak1964some; defazio2020mom; ziyin2020laprop); see Appendix 8. We summarize these formulations and their differences in Table˜1.
@@
Table 1: Summary of Nesterov, Schedule-Free, and GPA formulations. Here, $\mu$ is the momentum parameter, $\gamma^{(t)}$ is the learning rate, and $g(\cdot;\xi^{(t)})$ denotes the stochastic gradient. For a summary of Polyak momentum formulations, please refer to Appendix 8.
Formulation Update Equations Notes
[…]
@@ 2.2 Single-Worker DiLoCo and its Weaknesses
DiLoCo was originally introduced as a distributed algorithm for cross-datacenter training (diloco). The method computes multiple inner steps of a base optimizer on the inner weights, then applies Nesterov (Equation˜3) on the average pseudo-gradient, the difference between the previous and updated inner model weights, to update the outer weights. The inner weights are then reset to outer weights.
DiLoCo requires storing two additional optimizer states of the same shape as the model parameters: the momentum buffer $b^{(t)}$ and the current model parameters $x^{(t)}$ (also known as the outer weights). DiLoCo’s handling of fast inner weights and slow outer weights can be interpreted as a modified Lookahead method that applies Nesterov momentum to the outer weight updates (lookahead). The method was recently analyzed in khaled2025understanding, and demonstrated significant compute factor gains in the non-distributed setting in kallusky2025snoo. A simplified version of non-distributed or single-worker DiLoCo with $H$ inner steps of the base optimizer can be described as:
$p^{(t)}$ $=x^{(t)}-\operatorname{\texttt{BaseOptIteration}}(x^{(t)};\{\gamma^{(j)}\}_{j=1}^{H},H)$
$b^{(t)}$ $=\mu b^{(t-1)}+p^{(t)}$ (6)
$x^{(t+1)}$ $=x^{(t)}-\tilde{\gamma}[\mu b^{(t)}+p^{(t)}],$
@@
Weaknesses in DiLoCo’s hierarchical framework.
However, this two-level structure is undesirable. From an algorithmic perspective, one would prefer to average iterates on-the-fly, as opposed to averaging trajectories that implicitly contain multiple iterations of the base optimizer. From the users’ perspective, the two-level structure introduces an additional copy of the model weights required to compute the pseudo-gradient, and introduces additional hyperparameters to tune, e.g., the inner and outer learning rates, momentum, and number of inner steps. Lastly, from the distributed training perspective, DiLoCo couples the number of inner steps as a hyperparameter for both local SGD as well as for the modified Nesterov algorithm, causing the algorithm’s performance to counterintuitively improve as the number of inner steps increases. One would instead expect that communicating more often should always be beneficial. These challenges motivate the development of a new algorithm that removes the two-level structure while offering a separate hyperparameter that can smoothly average the observed iterates at every iteration.
@@ 2.3 Schedule-Free Learning
In parallel, Schedule-Free learning (SF) (schedule-free) was recently proposed as a wrapper to any base optimizer using a variant of the primal averaging formulation of Nesterov’s method (Equation˜4) for hyperparameter-free learning:
@@
Originally designed to eliminate the need for manually specified learning rate schedules, Schedule-Free has demonstrated the surprising ability to not only match, but even surpass the practical performance of the original base optimizer. This is done by decoupling the momentum hyperparameter used in the $x^{(t)}$ and $y^{(t)}$ sequences, unlike the standard primal averaging formulation of Nesterov (Equation˜4). Through the choice of $\mu$, the method interpolates between uniform Polyak-Ruppert averaging and stochastic primal averaging (ruppert; polyak; tao2018).
@@ 3 Generalized Primal Averaging (GPA)
By decoupling the constants for the model evaluation and gradient computation sequences in the primal averaging formulation of Nesterov’s method (Equation˜4) and leveraging the observation of using exponential moving averaging in lieu of uniform averaging in Schedule-Free (Equation˜7), we introduce the Generalized Primal Averaging (GPA) framework:
@@
Unlike the modern formulation of Nesterov momentum (Equation˜3) or DiLoCo (Equation˜6) built on (pseudo-)gradient averaging, GPA is defined based on the primal or iterate averaging framework. We argue that this provides a more meaningful characterization of the method. For example, the primal averaging interpretation naturally extends to other search directions by replacing $g(y^{(t)};\xi^{(t)})$ with the search direction $d^{(t)}$ evaluated at $y^{(t)}$. This extension is not intuitive from the gradient averaging perspective, as it would translate to averaging search directions (with potentially different, evolving preconditioners) in the momentum buffer.
Learning rate schedules. By replacing Polyak-Ruppert averaging with exponential moving averaging, GPA is not inherently schedule-free and requires the use of a learning rate schedule. To see why, observe that Polyak averaging places increasingly less weight $1/(t+1)$ on the most recent iterate $z^{(t+1)}$, which plays a similar role to learning rate scheduling (sandler2023trainingtrajectories; schedule-free). GPA instead places a constant weight $\mu_{x}$ on the most recent iterate $z^{(t+1)}$ by leveraging an exponential moving average, thereby requiring a learning rate schedule compared to Schedule-Free. This is reflected theoretically in their average-iterate versus last-iterate convergence properties.
@@
Memory-efficient implementation. When $\mu_{y}>0$, GPA can be implemented with only one extra copy of the model weights – specifically, by storing $y^{(t)}$ and reconstructing $x^{(t)}$ from $y^{(t)}$ and $z^{(t)}$ during evaluation – unlike DiLoCo, which demands more memory overhead due to the global parameter and momentum buffers. The pseudocode is provided in Appendix 9.
Other properties. Because $\mu_{x},\mu_{y}\in[0,1]$, GPA preserves modular norm bounds of the model parameters (bernstein2024modular). More details on this property is provided in Appendix 9. GPA can also be interpreted as a special case of the Triple Momentum framework (7967721), which utilizes three momentum terms to achieve tighter global linear convergence bounds for deterministic, smooth, strongly convex functions compared to Nesterov’s method.